# Quantum Neural Networks in Practice: A Comparative Study with Classical Models from Standard Data Sets to Industrial Images

### v1

@@ III.2 Methodology 2: Convolutional neural networks
As already mentioned when introducing the first methodology in the previous Sec. III.1, the employed dimensionality reduction is a mere necessity — but one that heavily alters the classification problem being solved in its second stage. For our second methodology, we therefore avoid such dimensionality reduction methods and directly classify the full image data. However, while this is feasible for fully-classical models, it is still not feasible for fully-quantum models. This is why we now resort to a comparison between fully-classical and quantum-classical hybrid models, although in a way that dimensionality reduction is not needed. The hybrid models considered for this methodology consist of hybridizations of convolutional neural networks with quantum neural networks used as analogs of their convolution filters, which in classical convolutional layers are simple matrices. This hybrid model structure was first explored in Ref. [21]. One should not confuse these models with the ones referred to as quantum convolutional neural networks (QCNNs) [31], which are essentially fully-quantum neural networks that progressively trace out qubits in the middle of the circuit using, for instance, mid-circuit measurements. The latter are responsible for preventing barren plateaus in these models [32] and also add non-linearity to the quantum operations. Since the hybrid models used are much more computationally expensive to train due to many applications of each quantum circuit, we opt to preselect a smaller set of combinations of hyperparameters instead of relying on the randomization approach used before. However, we still train each model with $10$ different sets of random initial parameters for $100$ epochs to ensure that the results are not biased by the choice of initial parameters.
@@ IV.1 Comparison of classical and quantum neural network performances
Figure 5: Performance comparison between classical and quantum neural networks in terms of the classification accuracy of the best random models, i.e., the connected lines of Fig. 4 (a) but for all data sets.
@@
Figure 4 (a) shows the classification results for each of the $50$ randomly generated classical neural networks in orange colors. For the simplest classification task with $d=2$, our randomization methodology generates no models that perform excellent, i.e., achieve averaged classification accuracies close to $100\%$. In fact, even the best models do not surpass classification accuracies of around $\sim 75\%$ while the worst ones are just slightly above $50\%$. There is, however, a notable difference depending on size of the training data set. Unless for the smallest one with $N=200$, the larger data sets yield at best classification accuracies at around $60\%$. Given the rather drastic data reduction down to $d=2$ features, the poor overall classification accuracies are no surprise and the slightly better performance in case of $N=200$ may be attributed to the data set’s size. Figure 1 (i) exemplarily shows the data points as well as the learned function that classifies them best for $d=2$ and $N=500$. The corresponding classical neural network architecture is illustrated directly below Fig. 1 (i). Upon moving towards data reductions up to $d=16$, we observe an increase in the best model’s averaged classification accuracies to roughly $80\%$, despite none of the individual runs with one set of initial training parameters ever reaches $100\%$. For $d=32$ and beyond, our randomization approach starts yielding models giving rise to averaged classification accuracies up to $100\%$, indicating that a reduction to $d=32$ seems to capture enough information to warrant excellent classification.
@@
In Fig. 4 (c)-(h), we additionally analyze the variances of the classification accuracies given the $10$ sets of random initial training parameters whose after-training performance on the validation data set defines the averaged classification accuracies shown in Fig. 4 (a). It can be seen that when comparing results of classical and quantum models with similar averaged classification accuracies, their corresponding variances differ slightly. In fact, the quantum models variances seems to be generally smaller than those of the classical models. This is especially evident in Fig. 4 (e). Thus, although the quantum neural networks perform similar at maximum but worse on average than their classical counterparts in terms of classification accuracies, they yield those results more consistently and do depend less on the initial starting parameters used for training. This may be regarded advantageous if good initial training parameters are not known in advance.
@@
Moreover, we generally observe an increase in the accuracy variance of classical models as we increase the number of data points in the data set. This can be due to the additional data points increasing the separation hardness between the classes and leading to more complex loss landscapes. While the same was not observed for quantum models, this effect is clearer with higher data dimensions where we lack information about the performance of the latter due the high computational cost. Therefore, a deeper analysis would be needed for a fair comparison but it might be an explanation for the surprisingly well performance of the quantum model with $N=200$ and $d=2$ in Fig. 4 (a).
@@
So far, we have discussed the classification results for the various versions of the third data set under a reduction via PCA. For the first and second data sets, cf. Secs. II.1 and II.2, as well as the third one with CAE reduction, we summarize their main results in Fig. 5 (a)-(d). The various panels show exclusively the maximally achieved classification accuracies, from among the $50$ randomly generated classical and quantum neural networks, obtained on the validation data set after training on the training data set. We deem this the most crucial performance indicator for comparing classical and quantum performance.
@@
For the third data set with CAE as the dimensionality reduction method, we give the results in Fig. 5 (d). In fact, they look very similar to those presented in Fig. 4 in almost every aspect. Most prominently, they also confirm that the best classical and quantum models perform equally well despite the overall difficulty to classify the strongly dimensionality-reduced industrial images.
@@
In Fig. 5 (a), the results for the first data set, i.e., the linear split of the hypercube, are shown. While each version of the data set, distinguished by the number of data points $N$, achieves almost $100\%$ classification accuracies for classical models and small hypercube dimensions $d$, their accuracies fall for larger $d$. This effect is even stronger for data sets with fewer data points, which is explainable given the hypercube’s growing size but constant number of sampling points from which to learn the function that splits it. In contrast, the best quantum models achieve classification accuracies of $\sim 90\%$ at maximum while their average and lowest performances (data not shown) show inferior performance compared to classical models as well. This observation of inferior performance is in agreement with Ref. [16], which observed that a linear split through a hypercube seems to be surprisingly difficult to classify for quantum models. Nevertheless, since it confirms former observations, we deem it a confirmation of our methodology.
@@
In Fig. 5 (b), the results for the images of handwritten digits, dimensionality-reduced by a CAE, are shown. While each version of the data set, distinguished by their reduced dimension $d$ and the number of data points $N$, achieves at least a classification accuracy of $>95\%$ for both classical and quantum models, the accuracies for the quantum models are still slightly lower. This becomes even more pronounced when using a PCA instead of a CAE for dimensionality reduction with the corresponding results shown in Fig. 5 (c). However, given the simplicity of the classification task that originates from the simplicity of the corresponding dimensionality reduced data, cf. Fig. 1 (a) and (b), it is not surprising, that both classical and quantum models solve the task well. This is in line with other studies targeting the same data set and which report similar classification accuracies [47, 48]. This serves as another confirmation of our methodology.
@@
At last, we want to exemplarily compare the learned functions of the best performing classical and quantum neural networks for all three data sets and both reduction methods to $d=2$, which is unfortunately the only dimension for which we can visualize them. The results are shown in Fig. 1 (f)-(j) and (k)-(o) for classical and quantum models, respectively, and the corresponding best performing model architectures are shown directly below or above the panels, respectively. We observe across all data sets that the quantum models tend to learn more complex functions than the classical ones despite some of the data’s simplicity, especially for the first and second data sets. Note that the classically learned functions do not change qualitatively or in simplicity even if replacing the linear activation function $\mathrm{ReLU}(x)=\max(0,x)$ with a non-linear one like $\tanh(x)$ (data not shown). Following from Ref. [49], this may help evaluate if the considered quantum neural networks have a poor inductive bias. Moreover, we also want to stress that we observe the same behavior of the variances as discussed in Fig. 4 (c)-(h). In fact, the effect of the quantum models having generally a lower variance with respect to the starting parameters is even stronger for the first and second data sets. Hence, it is an observation that holds for all examined data sets.
@@ IV.3 Analysis of correlations between performance and hyperparameters
Figure 7: Correlation analysis for the $50$ random quantum neural networks in terms of correlations between classification accuracy and various properties of the parametrized quantum circuit. The five rows correspond to the results obtained for the five data sets as indicated on the right. For simplicity and in order to have more data points for each data set, we merge all versions of each data set. To be precise, this implies that the presented data for the first data set contains the results gathered independently from training results with $N\in\{200,500,1000,2000\}$ data points. For the second and third data set, it implies joining results gathered independently by training them with $N\in\{200,500\}$. The colors distinguish the data dimensions. In panels (a)-(e),the correlation between the joint number of trainable parameters in the PQC $\mathsf{U}_{\mathrm{pqc}}(\boldsymbol{x}_{i},\boldsymbol{\phi},\boldsymbol{\theta})$, cf. Eqs. (5) and (6), as well as its measurement operator $\mathsf{O}(\boldsymbol{\omega})$, cf. Eqs. (8a) and (8), and the classification accuracy is analyzed. In panels (f)-(j), the correlation between the number $n_{\mathrm{layer},\mathrm{q}}$ of layers in the PQC and the classification accuracy is analyzed. While panels (k)-(o) analyze the correlation between the number $N_{2\mathrm{q}}$ of entangling gates and the classification accuracy, panels (p)-(t) analyze how their averaged gate entangling power, measured by Eq. (18), for the trained model is correlated with the accuracy. The linear lines in every panel — except panels (f)-(j) — indicate the general trend of the corresponding data. Note that we do not take models yielding exactly $50\%$ classification accuracy after training into account for calculating the trend lines as those models are intrinsically untrainable, due to bad luck during randomization, and would otherwise impact the trend for actually trainable models.
@@
While Sec. IV.2 shows that our best randomly generated quantum neural networks achieve similar performances than some tested literature models, we now analyze whether there are some common features which distinguish well and badly performing random models. To this end, we first investigate the correlation between the achieved classification accuracies and the total number of trainable parameters in the model, since the latter is often connected to a model’s expressivity and thus performance. Figure 7 (a) analyzes this correlation for the first data set and reveals that the classification accuracy of the hypercube data for $d=2$ does not depend much on the number of trainable parameters, as low and high accuracies can be observed for both low and high numbers of trainable parameters. This changes for higher dimensional hypercubes, which show an increasing correlation between accuracy and the number of trainable parameters when moving to $d=4$ and then further to $d=8$. This behavior is even stronger for the two versions of the third data set of industrial images, distinguished by the employed reduction method and shown in Fig. 7 (d) and (e). A possible explanation would be that due to the increased complexity of the data, cf. Fig. 1 (d) and (e), it takes more parameters to express a sufficiently complex function that is able to at least partially approximate the exact but unknown classification function $f:\mathbb{R}^{d}\rightarrow\{0,1\}$. This is also in line with the results for the second data set presented in Fig. 7 (b) and (c), where the number of trainable parameters seems to play the least important role compared to the other two data sets. This is explainable by the underlying data’s low complexity, cf. Fig. 1 (a) and (b). Interestingly, we find that the correlations in Fig. 7 (a), (d) and (e) originate mainly from correlations between the classification accuracy and the number of trainable parameter in the PQC $\mathsf{U}_{\mathrm{pqc}}(\boldsymbol{x}_{i},\boldsymbol{\phi},\boldsymbol{\theta})$ and not so much due to correlations with the number of trainable parameters in the measurement operator $\mathsf{O}(\boldsymbol{\omega})$. If explicitly calculating the correlations statistics with respect to either of the two (data not shown), we find very similar statistics to that shown in Fig. 7 (a)-(e) for the PQC and observe no correlations for the measurement operator.
@@
Besides the correlations analyzed in Fig. 7, we also analyze the correlations between the classification accuracy and various other quantities (data not shown), for instance with respect to the number of qubits $n_{\mathrm{q}}$, the form of the measurement operator $\mathsf{O}(\boldsymbol{\omega})$, the form of the entangling structure or the employed feature encoding. We did not find any correlations for any of these quantities for any of the data sets. We therefore conclude that our correlation analysis in general does not yield any clear indicator or rule on how most of the varied hyperparameters of the random quantum neural networks should be chosen for ideal performance. The only deducible rule seems to be that having more trainable parameters in the quantum models seems to be almost always beneficial, at worse irrelevant, but never harmful.
However, we may deduce something when only analyzing the best-performing quantum models. Upon inspecting in detail the structure of the three best-performing quantum models for all combinations of data set, dimensionality reduction method and number of features, we report that only five out of these $90$ models were circuits without any entangling gates. Moreover, analyzing the cases where we have $d=8$ features per data point, only three out of the $30$ best models did not use all-to-all entanglement. For these specific cases in particular, the best-performing quantum model always used all-to-all entanglement.

### v2

@@ III.2 Methodology 2: Convolutional neural networks
As already mentioned when introducing the first methodology in the previous Sec. III.1, the employed dimensionality reduction is a mere necessity — but one that heavily alters the classification problem being solved in its second stage. For our second methodology, we therefore avoid such dimensionality reduction methods and directly classify the full image data. However, while this is feasible for fully-classical models, it is still not feasible for fully-quantum models. This is why we now resort to a comparison between fully-classical and quantum-classical hybrid models, although in a way that dimensionality reduction is not needed. The hybrid models considered for this methodology consist of hybridizations of convolutional neural networks with quantum neural networks used as analogs of their convolution filters, which in classical convolutional layers are simple matrices. This hybrid model structure was first explored in Ref. [24]. One should not confuse these models with the ones referred to as quantum convolutional neural networks (QCNNs) [34], which are essentially fully-quantum neural networks that progressively trace out qubits in the middle of the circuit using, for instance, mid-circuit measurements. The latter are responsible for preventing barren plateaus in these models [35] and also add non-linearity to the quantum operations. However, they also make QCNNs classically simulable [21]. Since the hybrid models used are much more computationally expensive to train due to many applications of each quantum circuit, we opt to preselect a smaller set of combinations of hyperparameters instead of relying on the randomization approach used before. However, we still train each model with $10$ different sets of random initial parameters for $100$ epochs to ensure that the results are not biased by the choice of initial parameters.
@@ IV.1 Comparison of classical and quantum neural network performances
Figure 5: Performance comparison between classical and quantum neural networks in terms of the classification accuracy of the best random models, i.e., the connected lines of Fig. 4 (a) but for all data sets. The shaded background indicates the standard deviation generated by the $10$ individual random initial parameter sets that make up the shown averaged best random model performance.
@@
Figure 4 (a) shows the classification results for each of the $50$ randomly generated classical neural networks in orange colors. For the simplest classification task with $d=2$, our randomization methodology generates no models that perform excellent, i.e., achieve averaged classification accuracies close to $100\%$. In fact, even the best models do not surpass classification accuracies of around $\sim 75\%$ while the worst ones are just slightly above $50\%$. There is, however, a notable difference depending on the size of the training data set. Unless for the smallest one with $N=200$, the larger data sets yield at best classification accuracies at around $60\%$. Given the rather drastic data reduction down to $d=2$ features, the poor overall classification accuracies are no surprise and the slightly better performance in case of $N=200$ may be attributed to an artificially less complex task determined by the small validation data set. Figure 1 (i) exemplarily shows the data points as well as the learned function that classifies them best for $d=2$ and $N=500$. The corresponding classical neural network architecture is illustrated directly below Fig. 1 (i). Upon moving towards data reductions up to $d=16$, we observe an increase in the best model’s averaged classification accuracies to roughly $80\%$, despite none of the individual runs with one set of initial training parameters ever reaches $100\%$. For $d=32$ and beyond, our randomization approach starts yielding models giving rise to averaged classification accuracies up to $100\%$, indicating that a reduction to $d=32$ seems to capture enough information to warrant excellent classification.
@@
In Fig. 4 (c)-(h), we additionally analyze the variances of the classification accuracies given the $10$ sets of random initial training parameters whose after-training performance on the validation data set defines the averaged classification accuracies shown in Fig. 4 (a). It can be seen that when comparing results of classical and quantum models with similar averaged classification accuracies, their corresponding variances differ slightly. In fact, the quantum model variances seem to be generally smaller than those of the classical models. This is especially evident in Fig. 4 (e). Thus, although the quantum neural networks perform similar at maximum but worse on average than their classical counterparts in terms of classification accuracies, they yield those results more consistently and do depend less on the initial starting parameters used for training. While a deeper statistical analysis is needed to confirm the generality of this property, this can be advantageous if good initial training parameters are not known in advance.
@@
Moreover, we generally observe an increase in the accuracy variance of classical models as we increase the number of data points in the data set. This can be due to the additional data points increasing the separation hardness between the classes and leading to more complex loss landscapes. While the same was not observed for quantum models, this effect is clearer with higher data dimensions where we lack information about the performance of the latter due the high computational cost. Therefore, a deeper analysis would be needed for a fair comparison but it might be an explanation for the surprisingly good performance of the quantum model with $N=200$ and $d=2$ in Fig. 4 (a).
@@
So far, we have discussed the classification results for the various versions of the third data set under a reduction via PCA. For the first and second data sets, cf. Secs. II.1 and II.2, as well as the third one with CAE reduction, we summarize their main results in Fig. 5 (a)-(d). The various panels show exclusively the maximally achieved classification accuracies, from among the $50$ randomly generated classical and quantum neural networks, obtained on the validation data set after training on the training data set. We deem this the most crucial performance indicator for comparing classical and quantum performances.
@@
For the third data set with CAE as the dimensionality reduction method, we give the results in Fig. 5 (d). In fact, they look very similar to those presented in Fig. 4 in almost every aspect. Most prominently, they also confirm that the best classical and quantum models perform equally well despite the overall difficulty to classify the strongly dimensionality-reduced industrial images. All classical and quantum model performances lie furthermore within one standard deviation of each other, thereby statistically confirming that none outperforms the other.
@@
In Fig. 5 (a), the results for the first data set, i.e., the linear split of the hypercube, are shown. While each version of the data set, distinguished by the number of data points $N$, achieves almost $100\%$ classification accuracies for classical models and small hypercube dimensions $d$, their accuracies fall for larger $d$. This effect is even stronger for data sets with fewer data points, which is explainable given the hypercube’s growing size but constant number of sampling points from which to learn the function that splits it. In contrast, the best quantum models achieve classification accuracies of $\sim 90\%$ at maximum while their average and lowest performances (data not shown) show inferior performance compared to classical models as well. Moreover, not even the best models’ standard deviation reaches that of the classical models. This observation of inferior performance is in agreement with Ref. [19], which observed that a linear split through a hypercube seems to be surprisingly difficult to classify for quantum models. Nevertheless, by confirming former observations, we consider this a validation of our methodology.
@@
In Fig. 5 (b), the results for the images of handwritten digits, dimensionality-reduced by a CAE, are shown. While each version of the data set, distinguished by their reduced dimension $d$ and the number of data points $N$, achieves at least a classification accuracy of $>95\%$ for both classical and quantum models, the accuracies for the quantum models are still slightly lower. This becomes even more pronounced when using a PCA instead of a CAE for dimensionality reduction with the corresponding results shown in Fig. 5 (c), although their standard deviations stay in touch with those of the classical models. However, given the simplicity of the classification task that originates from the simplicity of the corresponding dimensionality reduced data, cf. Fig. 1 (a) and (b), it is not surprising, that both classical and quantum models solve the task well. This is in line with other studies targeting the same data set and which report similar classification accuracies [50, 51]. This serves as another validation of our methodology.
Figure 6: Overview of the variances as shown in Fig. 4 (c)-(h) but for all data sets and reduction methods as indicated on the right-hand side of each row. We only show it for those data dimensions $d$ for which we have results for both classical and quantum neural networks.
@@
At last, we want to exemplarily compare the learned functions of the best performing classical and quantum neural networks for all three data sets and both reduction methods to $d=2$, which is unfortunately the only dimension for which we can visualize them. The results are shown in Fig. 1 (f)-(j) and (k)-(o) for classical and quantum models, respectively, and the corresponding best performing model architectures are shown directly below or above the panels, respectively. We observe across all data sets that the quantum models tend to learn more complex functions than the classical ones despite some of the data’s simplicity, especially for the first and second data sets. Note that the classically learned functions do not change qualitatively or in simplicity even if replacing the linear activation function $\mathrm{ReLU}(x)=\max(0,x)$ with a non-linear one like $\tanh(x)$ (data not shown). Following from Ref. [52], this may help evaluate if the considered quantum neural networks have a poor inductive bias. Moreover, we also want to stress that we observe the same behavior of the variances as discussed in Fig. 4 (c)-(h). We provide an overview of the variances with respect to the other data sets and reduction methods in Fig. 6. In fact, the effect of the quantum models generally having lower variance with respect to the starting parameters is even stronger for the first and second data sets. Hence, this observation holds for all examined data sets.
@@ IV.3 Analysis of correlations between performance and hyperparameters
Figure 8: Correlation analysis for the $50$ random quantum neural networks in terms of correlations between classification accuracy and various properties of the parametrized quantum circuit. The five rows correspond to the results obtained for the five data sets as indicated on the right. For simplicity and in order to have more data points for each data set, we merge all versions of each data set. To be precise, this implies that the presented data for the first data set contains the results gathered independently from training results with $N\in\{200,500,1000,2000\}$ data points. For the second and third data set, it implies joining results gathered independently by training them with $N\in\{200,500\}$. The colors distinguish the data dimensions. In panels (a)-(e),the correlation between the joint number of trainable parameters in the PQC $\mathsf{U}_{\mathrm{pqc}}(\boldsymbol{x}_{i},\boldsymbol{\phi},\boldsymbol{\theta})$, cf. Eqs. (5) and (6), as well as its measurement operator $\mathsf{O}(\boldsymbol{\omega})$, cf. Eqs. (8a) and (8), and the classification accuracy is analyzed. In panels (f)-(j), the correlation between the number $n_{\mathrm{layer},\mathrm{q}}$ of layers in the PQC and the classification accuracy is analyzed. While panels (k)-(o) analyze the correlation between the number $N_{2\mathrm{q}}$ of entangling gates and the classification accuracy, panels (p)-(t) analyze how their averaged gate entangling power, measured by Eq. (18), for the trained model is correlated with the accuracy. The linear lines in every panel — except panels (f)-(j) — solely indicate the general trend of the corresponding data and should not be taken as a fit to the data. Note that we do not take models yielding exactly $50\%$ classification accuracy after training into account for calculating the trend lines as those models are intrinsically untrainable, due to bad luck during randomization, and would otherwise impact the trend for actually trainable models.
@@
While Sec. IV.2 shows that our best randomly generated quantum neural networks achieve similar performances than some tested literature models, we now analyze whether there are some common features which distinguish well and badly performing random models. To this end, we first investigate the correlation between the achieved classification accuracies and the total number of trainable parameters in the model, since the latter is often connected to a model’s expressivity and thus performance. Figure 8 (a) analyzes this correlation for the first data set and reveals that the classification accuracy of the hypercube data for $d=2$ does not depend much on the number of trainable parameters, as low and high accuracies can be observed for both low and high numbers of trainable parameters. This changes for higher dimensional hypercubes, which show an increasing correlation between accuracy and the number of trainable parameters when moving to $d=4$ and then further to $d=8$. This behavior is even stronger for the two versions of the third data set of industrial images, distinguished by the employed reduction method and shown in Fig. 8 (d) and (e). A possible explanation would be that due to the increased complexity of the data, cf. Fig. 1 (d) and (e), it takes more parameters to express a sufficiently complex function that is able to at least partially approximate the exact but unknown classification function $f:\mathbb{R}^{d}\rightarrow\{0,1\}$. This is also in line with the results for the second data set presented in Fig. 8 (b) and (c), where the number of trainable parameters seems to play the least important role compared to the other two data sets. This is explainable by the underlying data’s low complexity, cf. Fig. 1 (a) and (b). Interestingly, we find that the correlations in Fig. 8 (a), (d) and (e) originate mainly from correlations between the classification accuracy and the number of trainable parameters in the PQC $\mathsf{U}_{\mathrm{pqc}}(\boldsymbol{x}_{i},\boldsymbol{\phi},\boldsymbol{\theta})$ and not so much due to correlations with the number of trainable parameters in the measurement operator $\mathsf{O}(\boldsymbol{\omega})$. If explicitly calculating the correlations statistics with respect to either of the two (data not shown), we find very similar statistics to that shown in Fig. 8 (a)-(e) for the PQC and observe no correlations for the measurement operator.
@@
Besides the correlations analyzed in Fig. 8, we also analyze the correlations between the classification accuracy and various other quantities (data not shown), for instance with respect to the number of qubits $n_{\mathrm{q}}$, the form of the measurement operator $\mathsf{O}(\boldsymbol{\omega})$, the form of the entangling structure or the employed feature encoding. We did not find any correlations for any of these quantities for any of the data sets. We therefore conclude that our correlation analysis in general does not yield any clear indicator or rule on how most of the varied hyperparameters of the random quantum neural networks should be chosen for ideal performance. The only deducible rule seems to be that having more trainable parameters in the quantum models is almost always beneficial at the tested problem scale, and irrelevant at worse.
However, we may deduce something when analyzing only the best-performing quantum models. Upon inspecting in detail the structure of the three best-performing quantum models for all combinations of data set, dimensionality reduction method and number of features, we report that only five out of these $90$ models were circuits without any entangling gates. Moreover, analyzing the cases where we have $d=8$ features per data point, only three out of the $30$ best models did not use all-to-all entanglement. For these specific cases in particular, the best-performing quantum model always used all-to-all entanglement.